# The algebraic and geometric classification of

### v1

@@ 2.5. The geometric classification of compatible pre-Lie algebras
The rest of the non-degenerations between indicated algebras follows from Remark 18.
Thanks to Theorem 21, we have that $\mathcal{C}_{33},$ $\mathcal{C}_{09}^{\alpha,\beta},$ $\mathcal{C}_{22}^{\alpha,\beta},$ $\mathcal{C}_{39}^{\alpha,\beta},$ $\mathcal{C}_{31}^{\alpha,\beta,\gamma},$ and $\mathcal{C}_{38}^{\alpha,\beta,\gamma}$ give irreducible components in the variety of $2$-dimensional compatible Novikov algebras. Hence, each $2$-dimensional compatible Novikov algebras is on the orbit closure of these algebras. The rest of the necessary degenerations are given below.
$\mathcal{C}_{24}^{1,1+t^{-1},-1+t^{-1}}\xrightarrow{(te_{1}+(-1+t)e_{2},\ t^{2}e_{1}+te_{2})}\mathcal{C}_{02},$ $\mathcal{C}_{24}^{1,t^{-1},\alpha t^{-1}}\xrightarrow{(te_{1}+\frac{t^{2}}{1-\alpha}e_{2},\ t^{2}e_{1}+e_{2})}\mathcal{C}_{05}^{\alpha\neq 1},$ $\mathcal{C}_{40}^{t^{-1},0,\frac{1}{2t}}\xrightarrow{(te_{1},\ te_{2})}\mathcal{C}_{08},$ $\mathcal{C}_{11}^{\alpha,\alpha+t,}\xrightarrow{(e_{1}+\beta t^{-1}e_{2},\ e_{2})}\mathcal{C}_{12}^{\alpha,\beta},$ $\mathcal{C}_{24}^{-1+\alpha,t^{-1},\alpha t^{-1}}\xrightarrow{(te_{1}-\frac{t^{2}}{\alpha-1}e_{2},\ t^{2}e_{1}-t^{3}e_{2})}\mathcal{C}_{16}^{\alpha\neq 1},$ $\mathcal{C}_{24}^{-2,\alpha+t^{-1},-\alpha+t^{-1}}\xrightarrow{(te_{1}-\alpha^{-1}te_{2},\ t^{2}e_{1}+2\alpha^{-1}t^{2}e_{2})}\mathcal{C}_{17}^{\alpha\neq 0},$ $\mathcal{C}_{40}^{-\alpha-2t^{-2},-6\alpha-6t^{-2},-\frac{4+\alpha t^{2}}{6t^{2}}}\xrightarrow{(te_{1}-3te_{2},\frac{t^{2}}{2}e_{1}-3t^{2}e_{2})}\mathcal{C}_{19}^{\alpha},$ $\mathcal{C}_{24}^{-1-t^{-2},-(\beta+t)t^{-2},(t-\alpha)t^{-2}}\xrightarrow{(-t^{3}e_{1}+\frac{t^{4}+t^{6}}{1+\alpha t-\beta t}e_{2},\ -t^{2}e_{1}+\frac{t^{5}}{\beta t-\alpha t-1}e_{2})}\mathcal{C}_{20}^{\alpha,\beta},$ $\mathcal{C}_{24}^{-1-t^{-2},-\beta t^{-2},-\alpha t^{-2}}\xrightarrow{(-t^{3}e_{1}+\frac{t^{3}+t^{5}}{\alpha-\beta}e_{2},\ -t^{2}e_{1}-\frac{t^{4}}{\alpha-\beta}e_{2})}\mathcal{C}_{21}^{\alpha,\beta},$ $\mathcal{C}_{27}^{\alpha-5t^{2},t^{-2}}\xrightarrow{(e_{1}+2t^{2}e_{2},\ te_{2})}\mathcal{C}_{26}^{\alpha},$ $\mathcal{C}_{24}^{\alpha,\beta,\gamma}\xrightarrow{(e_{1},\ t^{-1}e_{2})}\mathcal{C}_{25}^{\alpha,\beta,\gamma}.$

### v2

@@ 3.5 The geometric classification of compatible pre-Lie algebras
The rest of the non-degenerations between indicated algebras follows from Remark 3.6.
Thanks to Theorem 3.9, we have that $\mathcal{C}_{33}$, $\mathcal{C}_{09}^{\alpha,\beta}$, $\mathcal{C}_{22}^{\alpha,\beta}$, $\mathcal{C}_{39}^{\alpha,\beta}$, $\mathcal{C}_{31}^{\alpha,\beta,\gamma}$, and $\mathcal{C}_{38}^{\alpha,\beta,\gamma}$ give irreducible components in the variety of $2$-dimensional compatible Novikov algebras. Hence, each $2$-dimensional compatible Novikov algebras is on the orbit closure of these algebras. The rest of the necessary degenerations are given by
$\mathcal{C}_{24}^{1,1+t^{-1},-1+t^{-1}}\xrightarrow{(te_{1}+(-1+t)e_{2},t^{2}e_{1}+te_{2})}\mathcal{C}_{02},\quad\mathcal{C}_{24}^{1,t^{-1},\alpha t^{-1}}\xrightarrow{(te_{1}+\frac{t^{2}}{1-\alpha}e_{2},t^{2}e_{1}+e_{2})}\mathcal{C}_{05}^{\alpha\neq 1},$
$\mathcal{C}_{40}^{t^{-1},0,\frac{1}{2t}}\xrightarrow{(te_{1},te_{2})}\mathcal{C}_{08},\quad\mathcal{C}_{11}^{\alpha,\alpha+t,}\xrightarrow{(e_{1}+\beta t^{-1}e_{2},e_{2})}\mathcal{C}_{12}^{\alpha,\beta},$
$\mathcal{C}_{24}^{-1+\alpha,t^{-1},\alpha t^{-1}}\xrightarrow{(te_{1}-\frac{t^{2}}{\alpha-1}e_{2},t^{2}e_{1}-t^{3}e_{2})}\mathcal{C}_{16}^{\alpha\neq 1},$
$\mathcal{C}_{24}^{-2,\alpha+t^{-1},-\alpha+t^{-1}}\xrightarrow{(te_{1}-\alpha^{-1}te_{2},t^{2}e_{1}+2\alpha^{-1}t^{2}e_{2})}\mathcal{C}_{17}^{\alpha\neq 0},$
$\mathcal{C}_{40}^{-\alpha-2t^{-2},-6\alpha-6t^{-2},-\frac{4+\alpha t^{2}}{6t^{2}}}\xrightarrow{(te_{1}-3te_{2},\frac{t^{2}}{2}e_{1}-3t^{2}e_{2})}\mathcal{C}_{19}^{\alpha},$
$\mathcal{C}_{24}^{-1-t^{-2},-(\beta+t)t^{-2},(t-\alpha)t^{-2}}\xrightarrow{(-t^{3}e_{1}+\frac{t^{4}+t^{6}}{1+\alpha t-\beta t}e_{2},-t^{2}e_{1}+\frac{t^{5}}{\beta t-\alpha t-1}e_{2})}\mathcal{C}_{20}^{\alpha,\beta},$
$\mathcal{C}_{24}^{-1-t^{-2},-\beta t^{-2},-\alpha t^{-2}}\xrightarrow{(-t^{3}e_{1}+\frac{t^{3}+t^{5}}{\alpha-\beta}e_{2},-t^{2}e_{1}-\frac{t^{4}}{\alpha-\beta}e_{2})}\mathcal{C}_{21}^{\alpha,\beta},\quad\mathcal{C}_{27}^{\alpha-5t^{2},t^{-2}}\xrightarrow{(e_{1}+2t^{2}e_{2},te_{2})}\mathcal{C}_{26}^{\alpha},$
$\mathcal{C}_{24}^{\alpha,\beta,\gamma}\xrightarrow{(e_{1},t^{-1}e_{2})}\mathcal{C}_{25}^{\alpha,\beta,\gamma}.$
Acknowledgements
The authors thank Amir Fernández Ouaridi for sharing some useful software programs. The work is supported by FCT UIDB/00212/2020, UIDP/00212/2020 and 2022.14950.CBM (Programa PESSOA).